Gender and handedness (right-handed or left-handed) are recorded for a randomly-selected sample of adults. Of the 100 women in the sample, 92 women are right-handed. Of the 80 men in the sample, 70 men are right-handed.
a.   Write a two-way table of observed counts.
b.   Calculate the conditional distribution of handedness, given gender.
c.   Describe the conditional distribution above in a few sentences.
d.   Determine expected counts for all combinations of gender and handedness.
e.   Calculate the value of the chi-square test statistic.
f.   Now, show all the steps of the hypothesis test, making sure to write out the conclusion in non-technical language.
Answer #1

15.4 Sex (female or male) and handedness (right-handed or left- handed) are recorded for a randomly selected sample of adults. Of the 100 women in the sample, 92 women are right-handed. Of the 80 men in the sample, 70 men are right-handed. a.Write a two-wav table of observed counts Gender Male Female Total Handedness Right 70 92 162 Handedness 10 18 Total 80 100 180

b)

given Male

P(right)=70/80

P(left)=10/80

given female

P(right)=92/100

P(left)=8/100

c)

in genral conditional probability for right- handedness is more than hat of left

d) , e) and f)

c1 c2 total
r1 70 10 80
r2 92 8 100
total 162 18 180
Oi 70 10 92 8
Ei 72 8 90 10
TS
(O--Ei)^2/Ei 0.055556 0.5 0.044444 0.4 1
critical value 3.84
p-value 0.317310508

TS = 1

TS < critical value

we fail to reject the null hypothesis

there is not enough evidence to conclude that Gender and handedness are dependent